OpenStudy (anonymous):

please simplify i^37

OpenStudy (anonymous):

do you know what i^2 = ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the definition is \(\large i^2=-1 \)

OpenStudy (anonymous):

notice that i^4 = (i^2)^2=(-1)^2 so i^4 = ???

OpenStudy (anonymous):

\[i=\sqrt{-1}\]\[i ^{37}=i ^{36+1} = i \times i ^{36}=i \times (i ^{4})^{9}\]\[i ^{4}=\sqrt{-1}^{4}=(\sqrt{-1}^{2})^{2}=1^{2}=1\] Thus,\[i \times (i^{4})^{9}=i \times (1)^{9}=i \times 1=i\]

OpenStudy (anonymous):

so i?